ID |
Date |
Author |
Category |
Subject |
Text |
 |
154
|
Wed Jun 30 14:48:54 2021 |
Oliver | General | Data taking - and not... | We just started data taking. But a few minutes
later the main experiment needed to do some
adjustements to the beam. This interrupts |
|
153
|
Wed Jun 30 13:57:14 2021 |
Yuri, Sergey | Accelerator | sextupoles + scraper positions | After setting up on 70Se
Sextupoles were adjusted to optimize the
resolution of the 70Se peak. Set values are |
 |
152
|
Wed Jun 30 13:05:08 2021 |
Rui-Jiu Chen | Analysis | ID of 70Se | This is the latest ID of 70Se setting. |
    |
151
|
Wed Jun 30 13:04:40 2021 |
Shahab | Runs | 245 Amplifier on | turned on extra 20dB amplifier for the 245
Resonator around 12:32 |
|
150
|
Wed Jun 30 12:35:53 2021 |
Rui-Jiu Chen | Analysis | Change reference level to -20dBm. Turn of amplifier. | SC_2021-06-30_12_35_52
IQ_2021-06-30_12_35_43
Reference level=-20dBm |
|
149
|
Wed Jun 30 12:26:58 2021 |
Rui-Jiu Chen | Analysis | NTCAP start. | SC_2021-06-30_11-00-12
IQ_2021-06-30_11-00-12
reference level -80dBm.
|
|
148
|
Wed Jun 30 09:04:50 2021 |
Yuri, Markus | Accelerator | Status | 1) Over night the source was conservated.
It is being restarted now.
2) The approximated amount of Gas in this |
|
147
|
Wed Jun 30 08:59:37 2021 |
Yuri | Accelerator | 70Se setting | X(9Be)=1.815 g/cm2 (adjusted in primary beam
setting)
|
|
146
|
Wed Jun 30 02:12:46 2021 |
Shahab | Runs | still no beam in ESR: quadrupole magnets are down | still no beam in ESR: quadrupole magnets are
down |
|
145
|
Wed Jun 30 01:38:20 2021 |
Shahab | Runs | beam is back from Linac but not in the ESR | |
|
144
|
Wed Jun 30 00:20:10 2021 |
Rui-Jiu Chen | Analysis | ID of 70Se setting on May | This is an old spectrum which was recoded
on May. |
   |
143
|
Wed Jun 30 00:12:55 2021 |
Rui-Jiu Chen | Analysis | the ID of 72Ge setting | |
   |
142
|
Tue Jun 29 21:15:01 2021 |
Kanika, Shahab | Runs | lost beam problem Unilac | cooling water in UNILAC is out |
|
141
|
Tue Jun 29 21:12:15 2021 |
Kanika, Shahab | Runs | Moving the resonance frequency of the 245 MHz resonator | moving in the plungers:
before: 400
|
|
140
|
Tue Jun 29 21:10:30 2021 |
Shahab | Runs | 72-Ge setting | the new setting is roughly 400 khz lower in
frequency.
Previous beam time Ge was on 243.100 now |
|
139
|
Tue Jun 29 20:40:55 2021 |
Rui-Jiu Chen | DAQ | start NTCAP | Carrier Frequency:245MHz.
IQ rate 20MSamples/s
Reference level:-90 dBm
|
|
138
|
Tue Jun 29 20:13:58 2021 |
Helmut, Sergey, Yuri | Accelerator | change to predrilling an 72Ge | After optimisation of sextupoles we found
a scraper position which makes a narrow peak
and keeps 20% of the injected beam.
|
|
137
|
Tue Jun 29 10:32:10 2021 |
Sergey, Helmut | Accelerator | cooler curves for tuning of isochronicity | Variation of quadrupole E01QS03D to change
gamma_t in isoc. mode
measured with 78Kr at 368.5 MeV/u on 28.06.2021 |
 |
136
|
Mon Jun 28 18:32:49 2021 |
Ruijiu, Shahab | DAQ | trigger signal connected to NTCAP and analyzers | |
|
135
|
Fri Jun 25 21:47:24 2021 |
Yuri | Analysis | Identification | Attached are two files with a proposed particle
identification.
Different colors indicate different harmonic |
 |